OpenStudy (anonymous):

Solve for x in the proportion 3 over 8 = x over 12 x = 2 x = 4.5 x = 8.5 x = 9

OpenStudy (anonymous):

\(\frac{3}{8}=\frac{x}{12}\) means \(x=\frac{3\times 12}{8}\)